• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Abdurahim

    Log entry time 11:49:26 on June 3,2010

    Entry number 324249

    keyword=Compton Cavity is fine

    Compton cavity has ben checked out. Tweaked the PPLN temperature (decreased) and steering mirrors. Cavity is still fine and locking at around 2.5 kW at full laser power.
    Here are the parameters,

    === 06-03-2010-11-43 ===

    M1 (h, v) : 9695[c] 995[c]
    M2 (h, v) : 103310[c] -34095[c]
    M3 (v) : 280000[c]
    CavityStatus : Cavity is locked
    IPGPower (W): 5
    IPG Temp in (C): 42.4
    IPG Temp out (C): 33.25
    IPG Diode Current:2.308
    GreenPower (mW): 1506.38
    PPLN Temp (C) : 60.68
    Lightwave Temp(C):+47.298
    Lightwave Power: Pwr 84mW
    RampAmplitude(C): 0.08
    RampPeriod (s): 180
    Central Temp (C): 47.201
    Q1(x, y, s) : -111.736 588.822 -0.332052
    Q2(x, y, s) : 793.548 -32.8146 -0.668986
    PDR (V) : 0.231337
    PDT (V) : 0.735518
    Servo Gain : 120
    TrackingLevel (V):0.71
    Cavity Power (W): 2523.32
    S1 power (mW): 5.5127
    S2 power (mW): 369.404
    Polarization (%): 97.0592
    PolarizationState:LEFT

    ===



    FIGURE 1